The sum of two integers is 50. The larger is 2 more than three times the smaller. Write a system of equations to model this and then solve for the integers
Answer by stanbon(75887)

The sum of two integers is 50. The larger is 2 more than three times the smaller. Write a system of equations to model this and then solve for the integers.
--------
Equations:
x + y = 50
x = 3y + 2
----------------
Substitute for "x" and solve for "y"::
3y+2 + y = 50
4y = 48
y = 12
-----
x = 50-y = 50-12 = 38
--------------------
Cheers,
Stan H.
------------